OpenStudy (anonymous):

Find the zeros of the polynomial function. f(x) = x3 + 3x2 - x - 3

OpenStudy (***[isuru]***):

usual steps.... factorize it... x^3 + 3x^2 - x - 3 = ( x - 1 ) ( x^2 + 4x +3 ) = ( x - 1) ( x^2 + x + 3x + 3 ) = ( x -1 )[ x( x+ 1) + 3(x + 1) ] = ( x - 1) ( x + 3) (x + 1)

OpenStudy (***[isuru]***):

so... if u equal it to zero.. u will get x = 1 or x = -3 or x = -1 which means the zeros are 1 , -3 , -1